OpenStudy (anonymous):

what is the solution to the system of equations represented by the equation y equals negative 2x + 7 & y equals one third X

OpenStudy (alexandervonhumboldt2):

y=-2x+7 y=1/3x

OpenStudy (alexandervonhumboldt2):

1/3x=-2x+7